OpenStudy (anonymous):

4r squared minus 4r plus 1

OpenStudy (mysangels):

simplify the exponents so what is 4x4? then subtract 4r

OpenStudy (a1234):

Is that 4r^2 or (4r)^2?

OpenStudy (anonymous):

4r^2 - 4r + 1 = 0 (2r - 1)(2r - 1) = 0 (since it is factorable) r = -.5 and r = -.5

OpenStudy (anonymous):

Therefore, one root.

OpenStudy (mysangels):

then you can't have a number and number with a variable can't combine so you have s simplified expression